DEHN SURGERY ON ARBORESENT KNOTS AND LINKS – A SURVEY

YING-QING WU

Abstract. This article is solicited by C. Adams for a special issue of Chaos, Solitons and Fractals devoted to knot theory and its applications. We present some recent results about Dehn surgeries on arborescent knots and links.